Carter, Inc. produces two different products, Product A and Product B. Carter uses a traditional volume-based costing system in which direct labor hours are the allocation base. Carter is considering switching to an ABC system by splitting its manufacturing overhead cost of $783,000 across three activities: Design, Production, and Inspection. Under the traditional volume-based costing system, the predetermined overhead rate is $2.61/direct labor hour. Under the ABC system, the rate for each activity and usage of the activity drivers are as follows:

Activity Rate Usage by Product A Usage by Product B
Design (Engineering Hours) $ 500 /hour 100 300
Production (Direct Labor Hours) $ 1.25 100,000 200,000
Inspection (Batches) $ 520 300 100


Required:

a. Calculate the indirect manufacturing costs assigned to Product A under the traditional costing system.




b. Calculate the indirect manufacturing costs assigned to Product B under the traditional costing system.



c. Calculate the indirect manufacturing costs assigned to Product A under the ABC system.



d. Calculate the indirect manufacturing costs assigned to Product B under the ABC system.



e. Which product is under-costed and which is over-costed under the volume-based cost system compared to ABC?
Answer #1

a. Indirect manufacturing costs assigned to Product A under the traditional costing system = 100,000*2.61 = 261,000

b. Indirect manufacturing costs assigned to Product B under the traditional costing system = 200,000*2.61 = 522,000

c. Indirect manufacturing costs assigned to Product A under the ABC system = (500*100)+ (100,000*1.25)+ (300*520)

= 331,000

d. Indirect manufacturing costs assigned to Product A under the ABC system = (500*300)+ (200,000*1.25)+ (100*520)

= 452000

e. Product A is under costed and product B is overcosted.